# RsGCN: Rescaling Enhances Generalization of GCNs for Solving Scalable Traveling Salesman Problems

### v3

@@ 4.1.3 Architecture of GCN
Due to space limitations, this section has been placed in Appendix B.
@@ 4.2 Training Strategy
Since RsGCN acquires powerful generalization capability with the Rescaling Mechanism, a mixed-scale dataset is used to train RsGCN. In detail, the mixed-scale dataset includes one million TSP instances, in which the number of nodes varies in 20, 30, 50, and 100, and the proportion of these four scales is 1:2:3:4. On the one hand, the mixed-scale dataset effectively helps GCNs learn universal patterns across scalable TSPs, avoiding the pitfalls of narrow patterns specific to a single scale. On the other hand, the maximum scale in the mixed-scale dataset is only 100 nodes, and RsGCN can be efficiently generalized to 10K-node instances without any fine-tuning, which can significantly reduce the training cost.
@@
The single-label cross-entropy with Softmax is widely used in the loss function of neural TSP solvers. However, previous work does not consider that each node is connected to exactly two other nodes in TSP tours. Accordingly, we employ a double-label binary cross-entropy with Sigmoid as the loss function. In detail, the two adjacent nodes of $x_{i}$ in the optimal tour $\hat{\Pi}$ serve as the two labels. The loss function $\mathcal{L}$ for $n$-node instances is presented as follows:
$\mathcal{L}$ $=-\frac{\sum_{x_{i}\in\mathcal{X}_{n}}\sum_{x_{j}\in\mathcal{X}_{n}}\mathcal{Z}_{i,j}}{n}$ (4)
$\mathcal{Z}_{i,j}$ =$=\begin{cases}\log(H_{i,j}+\varepsilon),&\text{if }x_{j}\text{ isadjacent to }x_{i}\text{ in }\hat{\Pi}\\ \log(1-H_{i,j}+\varepsilon),&\text{others}\end{cases}$ (5)
where the division by $n$ in Equation 4 aims to balance the loss for scalable TSPs, helping prevent the gradient updates from being biased towards larger-scale instances. In Equation 5, $H_{i,j}$ is the probability between $x_{i}$ and $x_{j}$ in the heatmap normalized by the Sigmoid function. $\varepsilon$ is a very small value used to avoid taking the logarithm of zero.
4.3 Efficient Post-Search – Re2Opt
An efficient post-search algorithm is essential to fully utilize heatmaps. Inspired by the k-Opt operations in MCTS, we propose a more efficient and robust search algorithm termed Reconstruction-2Opt (Re2Opt). The overall procedure of Re2Opt is shown in Figure 1, which first conducts State Initialization to obtain an initial tour and then iteratively conducts Reconstruction and 2Opt and Weight Enhancement for further optimization until the time budget is exhausted.
@@ 4.3.1 State Initialization
The initial tour is constructed by greedy selection. First, a node $x_{i}$ is randomly selected from $\mathcal{X}_{n}$. Second, a candidate set of the next node for $x_{i}$ is constructed, denoted by $\mathcal{C}_{i}$. In detail, $\mathcal{C}_{i}$ consists of the top $k_{2}$ hottest nodes for $x_{i}$, i.e., the top $k_{2}$ nodes with the highest probability between $x_{i}$ according to the generated heatmap. Third, the hottest node in $\mathcal{C}_{i}$ that has not been traversed yet is selected as the next node. If all nodes in $\mathcal{C}_{i}$ have been traversed but the TSP tour is still incomplete, the nearest untraversed node to $x_{i}$ is selected as the next node. Then, the edge between $x_{i}$ and the next node is adopted, and $x_{i}$ is updated to the selected next node. The second and third steps are conducted iteratively until all nodes are traversed. Finally, the initial tour is obtained.
Figure 3: An example of reconstruction in Re2Opt.
@@ 4.3.2 Reconstruction
Reconstruction aims to enhance the search diversity and escape from local optima. First, a node is randomly selected as the split node. There are two edges directly connected to the split node in the tour, and the one with a smaller weight is selected and removed to form an acyclic path. Herein, the weights of edges are stored in $\boldsymbol{Q}$, where $\boldsymbol{Q}_{i,j}$ represents the importance of the edge $x_{i}$-$x_{j}$. $\boldsymbol{Q}$ is initialized as an all-zero symmetric matrix and is updated during the 2Opt and Weight Enhancement. Then, taking the split node as the start node, a reconstruction action is conducted on the current acyclic path. In detail, a target node is selected from non-adjacent nodes of the start node based on the probability distribution $t_{i,j}=\frac{\boldsymbol{Q}_{i,j}}{\sum_{x_{h}\in\mathcal{C}_{i}}\boldsymbol{Q}_{i,h}}$, where $t_{i,j}$ represents the probability of $x_{j}$ being selected as the target node with $x_{i}$ as the start node. Finally, the acyclic path is updated by removing a directly connected edge of the target node and adding a new edge between the start node and the target node. Note that after a reconstruction action, the start node is also updated, and multiple reconstruction actions are conducted in a reconstruction process.
@@
4.3.3 2Opt and Weight Enhancement
After the reconstruction, a tour is constructed by adding the last edge to the reconstructed acyclic path. Subsequently, 2Opt is employed to further optimize the tour, and the search space of 2Opt is also confined to the candidate set of each node, reducing the time complexity from $O(n^{2})$ to $O(nk_{2})$. In a swap operation in 2Opt, two selected edges $x_{i_{1}}$-$x_{j_{1}}$ and $x_{i_{2}}$-$x_{j_{2}}$ are transformed into $x_{i_{1}}$-$x_{j_{2}}$ and $x_{i_{2}}$-$x_{j_{1}}$. For every swap that obtains improvement to the tour, we update $\boldsymbol{Q}$ to enhance the weight of transformed edges as:
$\boldsymbol{Q}_{i,j}\leftarrow\boldsymbol{Q}_{i,j}+\exp{(-\frac{L(\Pi^{new})}{L(\Pi^{pre})})},\ \forall(i,j)\in\{(i_{1},j_{2}),(i_{2},j_{1})\}$ (6)
where $\Pi^{pre}$ is the tour before the swap and $\Pi^{new}$ is the improved tour after the swap. The weight update will guide the next reconstruction, helping the reconstructed path escape local optima while preserving the optimal subpaths. Ultimately, it facilitates the attainment of a better tour in subsequent 2Opt optimization.
@@ 5.1 Datasets
Let TSP-$n$ denote $n$-node TSP instances in uniform distribution. Our mixed-scale training set, abbreviated as TSP-Mix, is derived from a portion of the training data used by Joshi et al. [12] and the detailed features are shown in Section 4.2. As a common setting, the test sets TSP-20/50/100/200/500/1K/10K are taken from part of those used by Fu et al. [6], while TSP-2K/5K are the same as those used by Pan et al. [24]. In detail, the test sets consist of 1024 instances each for TSP-20/50/100; 128 instances each for TSP-200/500/1K; and 16 instances each for TSP-2K/5K/10K. For real-world instances, 78 instances with the number of nodes below 20K are selected from TSPLIB [26].
@@ 5.2 Training
In the proposed RsGCN, we set the number of graph convolutional layers $l=6$, the feature dimension $h=128$, and the maximum number of adjacent nodes $k_{1}=\min(50,n)$ for $n$-node instances. The Adam optimizer [13] with a cosine annealing scheduler [21] is employed for training on TSP-Mix. We set $epochs=3$ and $batch\ size=32$, with the learning rate decaying cosinely from 5e-4 to 0. All experiments are conducted on an NVIDIA H20 (96 GB) GPU and an AMD EPYC 9654 (96-Core @ 2.40GHz) CPU. In our configuration, RsGCN only requires approximately 10 minutes per training epoch.
@@ 5.3 Baselines
Our proposed solver is compared with: (1) Classical Solvers/Algorithms: Concorde (Applegate et al. 2006), LKH-3 (Helsgaun 2017), 2Opt (Croes 1958); (2) Learning-Based Constructive Solvers: H-TSP (Pan et al. 2023), LEHD (Luo et al. 2023), GLOP (Ye et al. 2024), DRHG (Li et al. 2025); (3) Rule-Based Constructive Solvers: Att-GCN (Fu et al. 2021), SoftDist (Xia et al. 2024), DIFUSCO (Sun et al. 2023), Fast T2T (Li et al. 2024).
All baselines are evaluated on the same test sets and hardware platform as ours. We set $trials=10\text{K}$ for LKH-3 and limit the runtime to a reasonable maximum for Concorde to prevent unacceptable runtime when solving large-scale TSPs. For fairness in dealing with uniform distribution instances, we strive to control the total runtime of LC and RC solvers to be equivalent to or longer than ours. Additional reproduction details on baselines are provided in Appendix C.
@@ 5.4 Testing
For Re2Opt, we set the maximum number of threads to 128 to fully utilize the CPU, set the runtime limit to $0.05n$ seconds, and randomly sample the maximum number of reconstruction actions $M\in[10,\min(40,n))$ for $n$-node instances. To account for distributional differences, we set the candidate size $k_{2}$ to 5 and 10 for uniform distribution and real-world instances, respectively.
Three metrics are employed for performance evaluation, including (1) the average TSP tour length; (2) the average optimality gap; (3) the total solution time across all instances for each scale. Herein, the optimality gap of an instance is calculated as $(L/\hat{L})-1$, where $\hat{L}$ denotes the optimal tour length solved by Concorde. We conduct repeated experiments with 5 random seeds, reporting the average metrics for uniform distribution instances and the best metrics for real-world instances in TSPLIB.
@@ 5.5 Main Results
Tables 2 and 2, 3 (Table 3 on small-scale TSPs is shown in Appendix D) demonstrate RsGCN’s powerful generalization and state-of-the-art performance. With only 3 epochs of training on TSP-Mix containing instances with up to 100 nodes, RsGCN with Re2Opt as the post-search algorithm can be generalized to 10K-node instances. The proposed Re2Opt can also obtain competitive results without the guidance of RsGCN, validating the efficiency of its search framework. In addition, Figure 6 shows that RsGCN requires the fewest learnable parameters among neural baselines, and RsGCN also requires the fewest training epochs, as shown in Appendix G.
Table 1: Comparison results on TSP-1K/2K/5K/10K. G denotes greedy search. “—” indicates absence of specific reproduction code or available heatmaps. SoftDist generates heatmaps based on distance transform without neural networks. Re2Opt here employs the 5-nearest neighbors as candidates, substituting for 5-hottest neighbors generated by RsGCN to comparatively evaluate RsGCN’s guidance effect on Re2Opt. The two-stage solution time is connected with +. The first-stage times of Att-GCN and DIFUSCO are taken from their publications due to reproducibility issues. Except for Concorde and LKH-3, we mark the smallest and second smallest gaps in red and blue, respectively.
Solver Type TSP-1K TSP-2K TSP-5K TSP-10K
@@
Concorde Exact 23.1218 0.0000 1.75m 32.4932 0.0000 1.98m 51.0595 0.0000 5.00m 71.9782 0.0000 10.33m
LKH-3 Heuristic 23.1684 0.2017 17.96s 32.6478 0.4756 29.75s 51.4540 0.7727 2.06m 72.7601 1.0864 7.06m
2Opt Heuristic 24.7392 6.9961 0.12s 34.9866 7.6729 0.49s 55.1501 8.0115 7.87s 77.8939 8.2188 51.88s
H-TSP LC 24.6679 6.6868 36.97s 34.8984 7.4022 9.42s 55.0178 7.7523 20.08s 77.7446 8.0113 38.57s
LEHD LC 23.6907 2.4603 1.48m 34.2119 5.2892 2.15m 59.6522 16.8290 9.59m 90.5505 25.8026 54.45m
GLOP LC 23.8377 3.0962 1.51m 33.6595 3.5893 1.80m 53.1567 4.1073 4.56m 75.0439 4.2592 9.01m
DRHG LC 23.3217 0.8648 1.50m 32.8920 1.2274 2.01m 51.6350 1.1271 5.00m 72.8973 1.2770 9.00m
Att-GCN RC+MCTS 23.6454 2.2650 43.94s + 1.71m 74.3267 3.2627 4.16m + 17.38m
SoftDist MCTS 23.6622 2.3378 0.14s + 1.81m 33.3675 2.6909 0.12s + 3.39m 52.6930 3.1994 0.12s + 8.66m 74.1944 3.0791 0.14s + 17.03m
DIFUSCO RC+MCTS 23.4243 1.3083 0.12s + 1.71m 73.8913 2.6579 0.14s + 17.55m
Fast T2T RC+G+2Opt 23.3122 0.8236 2.95m 32.7940 0.9258 2.03m 51.7409 1.3345 4.35m 72.9450 1.3432 17.00m
Re2Opt Heuristic 23.2756 0.6657 52.38s 32.7452 0.7757 1.68m 51.5199 0.9018 4.18m 72.8742 1.2448 8.37m
[…]
Table 2: Average gaps(%) across different size intervals on TSPLIB instances. The results of other baselines are taken from previous works [19, 17]. Configuration settings are provided below solvers in the header. Complete results on each TSPLIB instance can be found in Appendix E.
Size DIFUSCO T2T Fast T2T BQ LEHD GLOP DRHG Re2Opt RsGCN

### v4

@@ 4.1.3 Architecture of GCN
The global node coordinates and the rescaled subgraph edge lengths serve as inputs to RsGCN, providing both global and local perspectives. Due to space limitations, this section has been placed in Appendix B.
@@ 4.2 Training Strategy
A mixed-scale dataset is used to train RsGCN. In detail, the mixed-scale dataset includes one million TSP instances, in which the number of nodes varies in 20, 30, 50, and 100, and the proportion of these four scales is 1:2:3:4. On the one hand, the mixed-scale dataset effectively helps RsGCN learn universal patterns across different scales, avoiding the pitfalls of narrow patterns specific to a single scale. On the other hand, the maximum scale in the mixed-scale dataset is only 100 nodes, and RsGCN can be efficiently generalized to 10K-node instances without any fine-tuning, which can significantly reduce the training cost.
@@
According to the undirected nature of the symmetric TSP, we employ a double-label binary cross-entropy as the loss function. In detail, the two adjacent nodes of $x_{i}$ in the optimal tour $\hat{\Pi}$ serve as the two labels. The loss function $\mathcal{L}$ for $n$-node instances is presented as follows:
$\mathcal{L}$ $=-\frac{\sum_{x_{i}\in\mathcal{X}_{n}}\sum_{x_{j}\in\mathcal{X}_{n}}\mathcal{Z}_{i,j}}{n},$ (4)
$\mathcal{Z}_{i,j}$ $=\begin{cases}\log(\boldsymbol{H}_{i,j}+\varepsilon),&\text{if }x_{j}\text{ is adjacent to }x_{i}\text{ in }\hat{\Pi},\\ \log(1-\boldsymbol{H}_{i,j}+\varepsilon),&\text{others},\end{cases}$ (5)
where the division by $n$ in Equation 4 aims to balance the loss on different scales, helping prevent the gradient updates from being biased towards larger-scale instances. In Equation 5, $\boldsymbol{H}_{i,j}$ represents the sigmoid-transformed probability of the edge $x_{i}\rightarrow x_{j}$. $\varepsilon$ is a very small value used to avoid taking the logarithm of zero.
4.3 Reconstruction-Based Search – RBS
Inspired by the k-Opt operations in MCTS (Fu et al., 2021), we propose a post-search algorithm termed Reconstruction-Based Search (RBS). RBS employs a candidate-based search strategy, enabling its search results to clearly reflect the quality of the heatmaps, thereby facilitating the evaluation of the generalization ability of GCNs. The overall procedure of RBS is shown in Figure 1, which first conducts State Initialization to obtain an initial tour and then iteratively conducts Reconstruction and Optimization and Edge Enhancement for further optimization until the time budget expires.
@@ 4.3.1 State Initialization
Let $k_{2}$ denote the size of the candidate set, and $\mathcal{C}_{i}$ denote the candidate set of $x_{i}$. Moreover, $\mathcal{C}_{i}$ consists of the top $k_{2}$ hottest nodes for $x_{i}$, i.e., the top $k_{2}$ nodes with the highest probability according to the generated heatmap $\boldsymbol{H}_{i}$. The initial tour is constructed by greedy selection. First, the start node $x_{i}$ is randomly selected from $\mathcal{X}_{n}$. Second, the hottest node in $\mathcal{C}_{i}$ that has not been traversed yet is selected as the next node. If all nodes in $\mathcal{C}_{i}$ have been traversed but the TSP tour is still incomplete, the nearest untraversed node is selected as the next node. The second step is conducted iteratively until all nodes are traversed.
Figure 3: An example of reconstruction in RBS.
@@ 4.3.2 Reconstruction
Reconstruction aims to enhance the search diversity and escape from local optima. First, a node is randomly selected as the split node. There are two edges directly connected to the split node in the tour, and the one with a smaller weight is selected and removed to form an acyclic path. Herein, the weights of edges are stored in $\boldsymbol{Q}$, where $\boldsymbol{Q}_{i,j}$ represents the importance of the edge $x_{i}$-$x_{j}$. $\boldsymbol{Q}$ is initialized as an all-zero symmetric matrix and is updated during the Optimization and Edge Enhancement. Then, taking the split node as the start node, a reconstruction action is conducted on the current acyclic path. In detail, a target node is selected from non-adjacent nodes of the start node based on the probability distribution $t_{i,j}=\frac{\boldsymbol{Q}_{i,j}+\varepsilon}{\sum_{x_{h}\in\mathcal{C}_{i}}(\boldsymbol{Q}_{i,h}+\varepsilon)}$, where $\varepsilon$ denotes a small number to prevent division by zero; $t_{i,j}$ represents the probability of $x_{j}$ being selected as the target node with $x_{i}$ as the start node. Finally, the acyclic path is updated by removing an edge of the target node and adding a new edge between the start node and the target node. Note that after a reconstruction action, the start node is also updated, and multiple reconstruction actions are conducted in a reconstruction process.
@@
4.3.3 Optimization and Edge Enhancement
After the reconstruction, a tour is constructed by closing the reconstructed acyclic path. Subsequently, 2-Opt is employed to further optimize the tour, and the search space of 2-Opt is also confined to the candidate set of each node. In a 2-Opt swap, two selected edges $x_{i_{1}}$-$x_{j_{1}}$ and $x_{i_{2}}$-$x_{j_{2}}$ are transformed into $x_{i_{1}}$-$x_{j_{2}}$ and $x_{i_{2}}$-$x_{j_{1}}$. For every swap that obtains improvement to the tour, we update $\boldsymbol{Q}$ to enhance the weight of transformed edges as:
$\boldsymbol{Q}_{i,j}\leftarrow\boldsymbol{Q}_{i,j}+\exp{(-\frac{L(\Pi^{new})}{L(\Pi^{pre})})},\ \forall(i,j)\in\{(i_{1},j_{2}),(i_{2},j_{1})\},$ (6)
where $\Pi^{pre}$ is the tour before the swap and $\Pi^{new}$ is the improved tour after the swap. The weight update will guide the next reconstruction, helping the reconstructed path escape local optima while preserving the optimal subpaths. Ultimately, it facilitates the attainment of a better tour in subsequent optimization.
@@ 5.1 Datasets
Let TSP-$n$ denote $n$-node TSP instances in uniform distribution. Our mixed-scale training set, abbreviated as TSP-Mix, is derived from a portion of the training data used by Joshi et al. (2019) and the detailed features are shown in Section 4.2. As a common setting, the test sets TSP-20/50/100/200/500/1K/10K are taken from part of those used by Fu et al. (2021), while TSP-2K/5K are the same as those used by Pan et al. (2023). In detail, the test sets consist of 1024 instances each for TSP-20/50/100; 128 instances each for TSP-200/500/1K; and 16 instances each for TSP-2K/5K/10K. For real-world instances, 78 instances with the number of nodes below 20K are selected from TSPLIB (Reinelt, 1991).
@@ 5.2 Training
In RsGCN, we set the number of graph convolutional layers $l=6$, the feature dimension $h=128$, and the maximum number of adjacent nodes $k_{1}=\min(50,n)$ for $n$-node instances. The Adam optimizer (Kingma & Ba, 2015) with a cosine annealing scheduler (Loshchilov & Hutter, 2017) is employed for training on TSP-Mix. We set $epochs=3$ and batchsize=32batch\ size=32, with the learning rate decaying cosinely from 5e-4 to 0. All experiments are conducted on an NVIDIA H20 (96 GB) GPU and an AMD EPYC 9654 (96-Core @ 2.40GHz) CPU. In our configuration, RsGCN only requires approximately 10 minutes per training epoch.
@@ 5.3 Baselines
Our method is compared with: (1) Classical Solvers/Algorithms: Concorde (Applegate et al., 2006), LKH-3 (Helsgaun, 2017), 2-Opt (Croes, 1958); (2) Transformer-Based Methods: H-TSP (Pan et al., 2023), LEHD (Luo et al., 2023), GLOP (Ye et al., 2024), DRHG (Li et al., 2025); (3) GCN-Based Methods: Att-GCN (Fu et al., 2021), SoftDist (Xia et al., 2024), DIFUSCO (Sun & Yang, 2023), Fast T2T (Li et al., 2024).
All baselines are evaluated on the same test sets and hardware platform as ours. We set $trials=10\text{K}$ and enable specialspecial option for LKH-3; limit the runtime to a reasonable maximum for Concorde to prevent unacceptable runtime when solving large-scale TSPs. For fairness in dealing with uniform-distribution instances, we strive to control the total runtime of neural methods to be equivalent to or longer than ours. Additional reproduction details on baselines are provided in Appendix C.
@@ 5.4 Testing
For RBS, we set the maximum number of threads to 128 to fully utilize the CPU, set the runtime limit to $0.05n$ seconds, and randomly sample the maximum number of reconstruction actions $M\in[10,\min(40,n))$ for $n$-node instances. To account for distributional differences, we set the candidate size $k_{2}$ to 5 and 10 for uniform-distribution and real-world instances, respectively.
Three metrics are employed for performance evaluation, including (1) the average TSP tour length; (2) the average optimality gap; (3) the total solution time across all instances for each scale. Herein, the optimality gap of an instance is calculated as $(L/\hat{L})-1$, where $\hat{L}$ denotes the optimal tour length solved by Concorde. We conduct repeated experiments with 5 random seeds, reporting the average metrics for uniform-distribution instances and the best metrics for real-world instances in TSPLIB.
@@ 5.5 Main Results
Tables 1 and 2 demonstrate RsGCN’s strong generalization and state-of-the-art performance on both uniform-distribution and real-world TSPs. With only 3 epochs of training on TSP-Mix containing instances with up to 100 nodes, the two-stage framework RsGCN + RBS can generalize to 10K-node instances. RBS can also obtain competitive results without the guidance of RsGCN, validating its efficiency. In addition, Figure 5 shows that RsGCN requires the fewest learnable parameters among neural baselines, and RsGCN also requires the fewest training epochs, as shown in Appendix F.
Table 1: Comparison results on uniform-distribution TSPs. RBS here employs the 5-nearest neighbors as candidates, substituting for 5-hottest neighbors generated by RsGCN to comparatively evaluate RsGCN’s guidance effect on RBS. The two-stage solution time is connected with +. The first-stage times of Att-GCN and DIFUSCO are taken from their publications due to reproducibility issues. Except for Concorde and LKH-3, we mark the smallest and second smallest gaps in red and blue, respectively.
Method TSP-20 TSP-50 TSP-100 TSP-200 TSP-500
Concorde 3.8403 0.0000 1.19s 5.6870 0.0000 3.35s 7.7560 0.0000 18.96s 10.7191 0.0000 7.44s 16.5464 0.0000 31.04s
LKH-3 3.8403 0.0000 28.32s 5.6870 0.0000 56.52s 7.7561 0.0003 55.33s 10.7193 0.0022 7.41s 16.5529 0.0391 10.12s
2-Opt 3.8459 0.1452 0.04s 5.7812 1.6490 0.07s 8.0124 3.3033 0.15s 11.2332 4.7946 0.08s 17.5688 6.1776 0.03s
H-TSP 10.8276 1.0122 6.09s 17.5962 6.3446 17.82s
LEHD 3.8403 0.0000 10.00s 5.6870 0.0005 25.00s 7.7572 0.0147 45.00s 10.7309 0.1103 15.00s 16.6603 0.6881 54.00s
GLOP 3.8406 0.0078 14.07s 5.6935 0.1143 29.42s 7.7839 0.3597 1.65m 10.7952 0.7099 19.86s 16.9089 2.1907 46.61s
DRHG 3.8403 0.0000 10.00s 5.6872 0.0039 25.00s 7.7581 0.0272 45.00s 10.7556 0.3410 15.00s 16.6448 0.5947 50.00s
Att-GCN 3.8403 0.0000 2.39s 5.6875 0.0090 15.91s 7.7571 0.0137 24.21s 10.7625 0.4035 20.62s 16.7996 1.5308 31.17s
+ MCTS + 10.14s + 23.16s + 46.13s + 12.99s + 51.53s
SoftDist 3.8745 0.8801 0.12s 5.7692 1.4495 0.12s 7.9553 2.5842 0.12s 10.8971 1.6647 0.12s 16.8161 1.6295 0.12s
+ MCTS + 9.82s + 23.93s + 45.89s + 13.39s + 51.53s
DIFUSCO 16.6378 0.5519 3.61m
+ MCTS + 51.40s
Fast T2T 3.8416 0.0348 53.00s 5.6876 0.0115 1.00m 7.7587 0.0337 1.23m 10.7445 0.2364 28.00s 16.6461 0.6024 37.00s
+ 2-Opt
RBS 3.8403 0.0000 10.31s 5.6874 0.0066 22.40s 7.7626 0.0854 42.14s 10.7473 0.2634 12.46s 16.6309 0.5108 27.15s
RsGCN 3.8303 0.0000 0.15s 5.6870 0.0000 0.20s 7.7570 0.0120 0.41s 10.7295 0.0966 0.12s 16.5941 0.2884 0.36s
+ RBS + 9.17s + 20.18s + 40.20s + 10.07s + 25.09s
Method TSP-1K TSP-2K TSP-5K TSP-10K
@@
Concorde 23.1218 0.0000 1.75m 32.4932 0.0000 1.98m 51.0595 0.0000 5.00m 71.9782 0.0000 10.33m
LKH-3 23.1684 0.2017 17.96s 32.6478 0.4756 29.75s 51.4540 0.7727 2.06m 72.7601 1.0864 7.06m
2-Opt 24.7392 6.9961 0.12s 34.9866 7.6729 0.49s 55.1501 8.0115 7.87s 77.8939 8.2188 51.88s
H-TSP 24.6679 6.6868 36.97s 34.8984 7.4022 9.42s 55.0178 7.7523 20.08s 77.7446 8.0113 38.57s
LEHD 23.6907 2.4603 1.48m 34.2119 5.2892 2.15m 59.6522 16.8290 9.59m 90.5505 25.8026 54.45m
GLOP 23.8377 3.0962 1.51m 33.6595 3.5893 1.80m 53.1567 4.1073 4.56m 75.0439 4.2592 9.01m
DRHG 23.3217 0.8648 1.50m 32.8920 1.2274 2.01m 51.6350 1.1271 5.00m 72.8973 1.2770 9.00m
Att-GCN 23.6454 2.2650 43.94s 74.3267 3.2627 4.16m
+ MCTS + 1.71m + 17.38m
SoftDist 23.6622 2.3378 0.14s 33.3675 2.6909 0.12s 52.6930 3.1994 0.12s 74.1944 3.0791 0.14s
+ MCTS + 1.81m + 3.39m + 8.66m + 17.03m
DIFUSCO 23.4243 1.3083 0.12s 73.8913 2.6579 0.14s
+ MCTS + 1.71m + 17.55m
Fast T2T 23.3122 0.8236 2.95m 32.7940 0.9258 2.03m 51.7409 1.3345 4.35m 72.9450 1.3432 17.00m
+ 2-Opt
RBS 23.2756 0.6657 52.38s 32.7452 0.7757 1.68m 51.5199 0.9018 4.18m 72.8742 1.2448 8.37m
RsGCN 23.2210 0.4293 0.94s 32.6784 0.5698 0.49s 51.3987 0.6645 2.83s 72.6335 0.9103 11.83s
+ RBS + 50.40s + 1.68m + 4.17m + 8.36m
Table 2: Average gaps(%) across different size intervals on TSPLIB instances. The results of other baselines are taken from previous works (Li et al., 2024; 2025). Configuration settings are provided below solvers in the header. Complete results on each TSPLIB instance can be found in Appendix D.
Size DIFUSCO T2T Fast T2T BQ LEHD GLOP DRHG RBS RsGCN